OpenStudy (anonymous):

What affect does h = 2 have on the function f(x) = log(x)? A. The graph moves 2 units to the right. B. The graph moves 2 units to up. C. The graph moves 2 units to the left. D. The graph is stretched by a factor of |2|.

OpenStudy (goformit100):

B. The graph moves 2 units to up.

OpenStudy (anonymous):

this works just like the quadratic form: y = a(x - h)² + k i.e. f(x) = alog(x - h) + k so h moves the curve left and right k moves the curve up and down